OpenStudy (anonymous):

2sqrt6 times sqrt10 I don't understand how to do it. Can you explain please : )

OpenStudy (anonymous):

If something is under a root that is raised to the same power you can multiply it for example \[\sqrt{a} \times \sqrt{b}\] \[\sqrt{ab}\]

OpenStudy (anonymous):